What volume of oxygen gas reacts with 5.00L of propane gas


1
Expert's answer
2021-07-05T03:24:03-0400

C3H8 + 5O2 = 3CO2 + 4H2O

n (C3H8) = V / 22.4 = 5.00 / 22.4 = 0.22 mol

According to the equation, n (O2) = 5 x n (C3H8) = 5 x 0.22 = 1.12 mol

V (O2) = n (O2) x 22.4 = 1.12 x 22.4 = 25 L